What is the expected value of perfect information of the following decision table? (hint: you probably need to calculate the expected value with perfect information and find the maximum EMVs of each option)

 

States of Nature

Alternatives

S1

S2

Option 1

200

300

Option 2

50

350

probability

.6

.4
Answer #1

Step - 1: The best outcome for the state of nature “S1" is “Option 1” with a payoff of 200. The best outcome for the state of nature “S2” is “Option 2” with a payoff of $350.

Expected value with perfect information EVwPI= (200)(0.6) + (350)(0.4) = 260.

Step - 2. We find EMV for each alternative:

Option 1 = 200 * 0.6 + 300 * 0.4 = 240

Option 2 = 50 * 0.6 + 350 * 0.4 = 170

The maximum EMV is $240 for Option 1, which is the expected outcome without perfect information.
Thus:
EVPI = EVwPI - Maximum EMV
= 260 - 240 = 20

Answer: 20